How many atoms are in each elemental sample?

3.5 mol Cu (Express your answer using two significant figures.)

7.5×10−3 mol C (Express your answer using two significant figures.)

15.2 mol Hg (Express your answer using three significant figures.)

0.190 mol Na (Express your answer using three significant figures.)
Answer #1

Answer:

Step 1:

Explanation:

Mole: The mass of substance containing the same number of fundamental units as there are atoms in exactly 12 g of 12C

No. of atoms = moles × Avogadro's number

Step 2: Solution

(a) No. of atoms = moles × Avogadro's number = 3.5 mol  × 6.022 × 1023 atoms/mol = 2.1× 1024 atoms

(b) No. of atoms = 7.5 × 10-3 mol  × 6.022 × 1023 atoms/mol = 4.5× 1021 atoms

(c) No. of atoms = 15.2 mol  × 6.022 × 1023 atoms/mol = 9.15× 1024 atoms

(d) No. of atoms = 0.190 mol  × 6.022 × 1023 atoms/mol = 1.14× 1023 atoms